# DDIL: Improved Diffusion Distillation with Imitating Learning

### v1

@@ 4.2 DDIL Integration
PD + DDIL: DDIL is integrated with progressive distillation using a DAgger-inspired approach Ross et al. (2011). Distillation is performed on mixed rollouts generated by alternating between the pre-trained and student diffusion models within each generation. A stateless DDIM solver facilitates this interleaved sampling process.
@@ 5 Experiments
Datasets and metrics: Following standard practice for evaluating text-to-image diffusion models Rombach et al. (2022); Meng et al. (2023), we evaluate our distilled models zero-shot on two public benchmarks: COCO 2017 (5K captions), and COCO 2014 Lin et al. (2014) (30K captions) validation sets. We use each caption to generate an image with a randomized seed and report CLIP score using OpenCLIP ViT-g/14 model Ilharco et al. (2021) to evaluate image-text alignment. We also report Fréchet Inception Distance (FID) Heusel et al. (2017) to estimate perceptual quality. To measure diversity of generation, we report $LPIPS_{Diversity}$, where for a given prompt we generate output for 10 different seeds and obtain pair-wise LPIPS score and finally average over 50 randomly sampled COCO 2017 prompts.
@@
Training: For all our experiments, we choose AdamW optimizer Loshchilov & Hutter (2017) with $051e-$ 05 learning rate with warmup and linear schedule on a batch size of 224 in case of progressive distillation, 360 in case of LCM and 7 in case of DMD2 on SSD1B. To optimize for GPU usage, we adopt gradient checkpoint and mixed-precision training. Please refer to Appendix A for additional training details.
@@
PD + DDIL:In case of progressive distillation, we train the model for 4k steps for $\epsilon$ to $v$ space conversion to perform step distillation in $v$ space Salimans & Ho (2021). Then we perform guidance conditioning following the same protocol as Meng et al. (2023) where we sample guidance scale $\omega\sim[2,14]$ and incorporate additional guidance embedding as in Rombach et al. (2022) followed by step distillation. Overall we train ${}10K$ steps to obtain guidance conditioned checkpoint SD$(gc)$. For progressive distillation, we start with a 32-step discretization assumption for the pre-trained diffusion model and perform $32\to 16$ step distillation for 5K iterations with 500 steps of warm-up. We progressively increase training compute or gradient steps as we go towards fewer iteration student. For $16\to 8$, we follow a similar protocol, but we further split each stage of training into two parts. First, we do distillation for 6K steps to obtain a checkpoint and resume with warmup and $051e-$ 05 learning rate for another 4K steps of training. For $8\to 4$ and $4\to 2$ distillation, we first distill model for 8K steps followed by another 6K steps.We split single stage of training into two parts to exactly match training protocol of Step Distillation and DDIL. See supplemental for additional details. We adopt timesteps for discretization from the default config of the DPM++ solver i.e., for 4-step models our timesteps are $\{999,749,500,250\}$.
@@
LCM + DDIL: We trained both LCM and DDIL models on the Common Caption dataset for 8,000 steps, using the SDv1.5 checkpoint and a batch size of 60 on 6 A100 GPUs. To enhance consistency distillation, we introduced backward trajectory sampling. Specifically, we randomly selected a number of inference steps (3, 4, or 5) and obtained samples at specific timesteps along the backward trajectory. This enabled us to incorporate consistency distillation loss feedback not only on forward diffused latents but also on these backward trajectory latents within our framework.
DMD2 + DDIL: In this work we consider distilling SSD1B checkpoint Gupta et al. (2024) with DMD2 for computational efficiency. To achieve stable training within the DMD2 framework, which utilizes a teacher model and a "fake" critic, we update the fake critic ten times for every update of the student model. We present two sets of results. The first serves as an oracle experiment, establishing an upper bound on performance. Here, we distill SSD1B into the DMD2 student model using a batch size of 64, 70,000 gradient updates, and SDXL critics. We then compare this oracle experiment with SSD1B checkpoints trained with a significantly smaller batch size of 7 and 30,000 gradient steps on a single A100 node. Reducing the batch size drastically leads to unstable training and poor performance. However, by employing reflected diffusion distillation, we achieve improved training stability and a significant boost in performance, both quantitatively and qualitatively. Further performance gains are observed when incorporating a mixed rollout setting of DDIL within DMD2, as demonstrated in the table 3
@@
Student Selection Prior: Our protocol for student selection in trajectory collection follows standard practice from imitation learning. Where early in training, student’s performance is bad and hence we prioritize sampling more from $p_{data}(x)$ but as training progress and student’s performance is good we want to obtain expert feedback on mistakes that student makes i.e., address co-variate shift caused by feedback and training, inference mismatch but still sample from $p_{data}(x)$ to preserve marginal data distribution.
@@
SnapFusion* Li et al. (2023b) 8 16 24.20 0.300 N/A
Step Distillation* Meng et al. (2023) 8 8 26.90 0.300 N/A
Step Distillation* Meng et al. (2023) 4 4 26.40 0.300 N/A
UFOGen* Xu et al. (2023b) 1 1 22.5 0.311 N/A
ImagineFlash* $\ddagger$ Kohler et al. (2024) 2 2 34.7 0.301 N/A
@@
+ Reflected 4 4 24.25 0.306 0.59
+ DDIL 4 4 23.44 0.308 0.59
+ Reflected + DDIL 4 4 22.86 0.309 0.59
@@
5.1 Text-guided Image Generation
We also demonstrate effectiveness of DDIL with distribution matching techniques which adopt multi-step student like in DMD2. When we apply DMD2 to SSD1B, we can observe that DDIL improves FID from $31.77\to 27.72$ and clip score from $0.320\to 0.326$ and HPSv2 score from $0.302\to 0.304$.
To minimize the overall computational burden of distillation, all experiments utilizing DMD2 employed a single A100 node with a batch size of 7. For a direct comparison with the original DMD2 configuration, an oracle experiment is conducted, distilling SSD1B with DMD2 across four A100 GPU nodes with a batch size of 64 and 70,000 gradient updates. We observed that training DMD2 with a smaller batch size exhibited instability, even with 10 fake critic updates per student update. However, integrating reflected diffusion and DDIL yielded improved stability and performance, producing qualitative results comparable to the larger-scale experiment.
By incorporating the DDIL framework and the reflected diffusion distillation formulation, we demonstrate enhanced training stability and achieve strong performance with DMD2 and LCM using significantly smaller batch sizes and fewer gradient updates.
Diversity vs. quality trade-off: Adversarial distillation methods Sauer et al. (2023; 2024) exhibit a decrease in generation diversity (measured by $LPIPS_{Diversity}$) compared to the baseline. Ss discussed in previous section most common objectives of distribution matching also have mode-seeking objects Xie et al. (2024); Yin et al. (2024). This highlights the quality-diversity trade-off often encountered in generative models when fine-tuning for human preferences or specific applications.
Table 3: Text guided image generation results on $512\times 512$ COCO 2017-5K validation set. These results are obtained by adopting latent consistency distillation retrained for SSD1B and incorporating DDIL within DMD2 setting. We integrate DDIL into the DMD2 framework by unrolling just the student model exactly like DMD2 but also unrolling teacher to corresponding noise level too to better capture underlying data distribution and align gradient fields of student model and teacher model.
Model Steps FID []delimited-[][$\downarrow$][ ↓ ] CLIP []delimited-[][$\uparrow$][ ↑ ] HPSV2 []delimited-[][$\uparrow$][ ↑ ] LPIPSDiversity []delimited-[][$\uparrow$][ ↑ ]
SSD1B (guidance = 8) 20 30.23 0.336 0.297 0.48
SSD1B-LCM 4 35.23 0.311 0.282 0.45
SSD1B-DMD2 (batch = 64,70k) 4 26.56 0.337 0.309 0.49
SSD1B-DMD2 (batch = 7,30k) 4 31.77 0.320 0.302 0.52
+ Reflected (batch = 7,30k) 4 29.32 0.323 0.302 0.51
+ Reflected + DDIL (batch = 7,30k) 4 27.72 0.326 0.304 0.51
6 Conclusion
This work introduces DDIL, a novel framework for distilling diffusion models that addresses the challenge of covariate shift while preserving the marginal data distribution. Integrating DDIL with established distillation techniques, including Progressive Distillation, Consistency Distillation (LCM), and Distribution Matching based Distillation (DMD2), consistently yields quantitative and qualitative improvements. Furthermore, we also show that integrating DDIL within the DMD2 framework enhances training stability, reduces required batch sizes, and improves computational efficiency demonstrating wider applicability and practical usefulness.
Appendix A Additional Training Details
In case of progressive distillation, we use DPM++ multi-step Lu et al. (2022) solver (2-step) for all evaluations except co-variate shift analysis, where a state-less solver like DDIM Zhang et al. (2022) will ease switching between two different models and corresponding reverse processes. We train our model on an internal text-to-image dataset and only perform distillation without updating the teacher model. In case of consistency distillation with LCM, we adopt the default LCMScheduler. In the case of DMD2, we follow the same settings as the original repository except that in our work, we focus on SSD1B Gupta et al. (2024) instead of SDXL to reduce computational requirements for training. By default, we adopt reflected diffusion for all three diffusion distillation techniques we considered in this work.
We follow similar protocol as progressive distillation for DDIL but as we now distill using on text-to-image data using forward process and interactively mixed unrolled trajectories, we have this additional sampling prior hyper parameters. For $16\to 8$ distillation $75\%$ we obtain $z_{t}$ using forward diffusion making this part of training exactly equivalent to step Distillation. For remaining $25\%$ we choose data from mixed interaction unrolled trajectories where for first 6K steps, student is only selected $15\%$ of time for $z_{t}\to z_{t-2k}$ transition for trajectory collection, where as for next 4K iterations we choose student $80\%$ of time and train $50\%$ of time on collected trajectories. Effectively for first part of training we have prior $x\sim p_{data}(x)$ as $0.75*0.85$ followed by $0.5*0.2$ for later part of training within a stage. And follow same sampling prior protocol for two parts of training for each stage of progressive distillation with same hyper-parameters as step distillation. Overall we need ${}40K$ updates to obtain a 4-step checkpoint or ${}55K$ updates to obtain 2-step checkpoint for step distillation or DDIL, as our sampling is parallelized across GPUs we observe $<5\%$ overhead for DDIL over step distillation, it takes 2 days on single node of 8 A100 GPUs to perform distillation.
Appendix B Additional Evaluation Results

### v2

@@ 4.2 DDIL Integration
PD + DDIL: DDIL is integrated with progressive distillation using a DAgger-inspired approach [30]. Distillation is performed on mixed rollouts generated by alternating between the pre-trained and student diffusion models within each generation. A stateless DDIM solver facilitates this interleaved sampling process.
Computational overhead: DDIL only requires maintaining a replay Buffer $D_{Dagger}$ which stores intermediate latents from interactive reverse process i.e., switching between teacher and student models using $\hat{x}_{t}$ (Algorithm 1 of supplementary). The buffer is updated periodically by replacing older data with new data and only needs to store the input latents for training. Empirically, when applying DDIL to PD (method with the most gradient updates), we observed less than a 5% increase in training time.
@@
DDIL integration does not incur significant additional compute as we collect sampled intermediate time-step latents into a dataset and updating this dataset at a significantly lower frequency compared to gradient feedback, exactly like DAgger and elaborated in context of progressive distillation in supplementary material.
@@ 5 Experiments
Datasets and metrics: Following standard practice for evaluating text-to-image diffusion models [29, 25], we evaluate our distilled models zero-shot on two public benchmarks: COCO 2017 (5K captions), and COCO 2014 [17] (30K captions) validation sets. We use each caption to generate an image with a randomized seed and report CLIP score using OpenCLIP ViT-g/14 model [9] to evaluate image-text alignment. We also report Fréchet Inception Distance (FID) [7] to estimate perceptual quality.
To measure fidelity and diversity of generations we report Density-Coverage, Precision-Recall [3, 13, 26] and also $LPIPS_{Diversity}$. Coverage improves upon the recall metric to better quantify diversity by building nearest neighbour manifolds around the real samples, instead of the fake samples, as they have less outliers. We use a more effective and robust DINOv2 [27] feature space to on 30k prompts of COCO2014 to report Density-Coverage, Precision-Recall metrics across all experiments. In case of $LPIPS_{Diversity}$, for a given prompt we generate output for 10 different seeds and obtain pair-wise LPIPS score and finally average over 50 randomly sampled COCO 2017 prompts.
@@
Training: For all our experiments, we choose AdamW optimizer [19] with $051e-$ 05 learning rate with warmup and linear schedule on a batch size of 224 in case of progressive distillation, 360 in case of LCM and 7 in case of DMD2 on SSD1B. To optimize for GPU usage, we adopt gradient checkpoint and mixed-precision training. Please refer to Appendix A for additional training details.
@@
PD + DDIL:In case of progressive distillation, we train the model for 4k steps for $\epsilon$ to $v$ space conversion to perform step distillation in $v$ space [32]. Then we perform guidance conditioning following the same protocol as [25] where we sample guidance scale $\omega\sim[2,14]$ and incorporate additional guidance embedding as in [29] followed by step distillation. Overall we train ${}10K$ steps to obtain guidance conditioned checkpoint SD$(gc)$. For progressive distillation, we start with a 32-step discretization assumption for the pre-trained diffusion model and perform $32\to 16$ step distillation for 5K iterations with 500 steps of warm-up. We progressively increase training compute or gradient steps as we go towards fewer iteration student. We share more details in Section 2 of Appendix.
@@
LCM + DDIL: We trained both LCM and DDIL models on the Common Caption dataset for 8,000 steps, using the SDv1.5 checkpoint and a batch size of 60 on 6 A100 GPUs. To enhance consistency distillation, we introduced backward trajectory sampling. Specifically, we randomly select a number of inference steps (3, 4, or 5) and obtained samples at specific timesteps along the backward trajectory. This enabled us to incorporate consistency distillation loss feedback not only on forward diffused latents but also on these backward trajectory latents within our framework.
DMD2 + DDIL: In this work we consider distilling SSD1B checkpoint [6] with DMD2 for computational efficiency. To achieve stable training within the DMD2 framework, which utilizes a teacher model and a ‘fake’ critic, we update the fake critic ten times for every update of the student model as we adopt ‘SSD1B’ as critic instead of SDXL and hence allow more updates.
We distill SSD1B with DND2 and also integrate DDIL trained with significantly small batch size of 7 and 40,000 gradient steps on a single A100 node. By adopting reflected diffusion distillation, we achieve improved training stability and a significant boost in performance, both quantitatively and qualitatively. Further performance gains are observed when incorporating a mixed rollout setting of DDIL within DMD2, as demonstrated in the table 3. We can observe consistent boost in both quality i.e., FID@30k of 19.41% and diversity i.e, coverage boost of 19.38% over baseline DMD2 when DDIL is integrated.
@@
Student Selection Prior: Our protocol for student selection in trajectory collection follows standard practice from imitation learning. Where early in training, student’s performance is bad and hence we prioritize sampling more from $p_{data}(x)$ but as training progress and student’s performance is good we want to obtain expert feedback on mistakes that student makes i.e., address co-variate shift caused by feedback and training-inference mismatch but still sample from $p_{data}(x)$ to preserve marginal data distribution.
Table 2: Text guided image generation results on MS-COCO 2014-30K validation set except for CLIP Score which is evaluated on COCO 2017-5K. Our 4-step Model demonstrates SOTA performance for checkpoints based on SD1.5 and improves on diversity
Model Steps NFEs FID@5K []delimited-[][$\downarrow$][ ↓ ] FID@30K []delimited-[][$\downarrow$][ ↓ ] CLIP []delimited-[][$\uparrow$][ ↑ ] Precision []delimited-[][$\uparrow$][ ↑ ] Recall []delimited-[][$\uparrow$][ ↑ ] Density []delimited-[][$\uparrow$][ ↑ ] Coverage []delimited-[][$\uparrow$][ ↑ ]
Progressive Distillation 4 4 23.34 14.72 0.302 0.595 0.605 0.301 0.397
+ DDIL 4 4 22.42 13.97 0.302 0.603 0.605 0.318 0.411
Progressive Distillation 2 2 26.43 16.46 0.288 0.392 0.543 0.153 0.202
+ DDIL 2 2 24.13 15.81 0.291 0.455 0.547 0.189 0.249
LCM 4 4 24.39 16.45 0.305 0.668 0.687 0.376 0.481
+ Reflected 4 4 24.25 16.44 0.306 0.670 0.687 0.378 0.478
+ DDIL 4 4 23.44 15.87 0.308 0.672 0.690 0.382 0.493
+ Reflected + DDIL 4 4 22.86 15.34 0.309 0.669 0.692 0.381 0.499
Improvement ($\mathbf{\Delta}$) 4 4 6.3% 6.7% 1.31% 0.14% 0.72% 1.32% 3.74%
SD [$v$] 32 64 22.50 13.48 0.321 - - - -
SD [$gc$, Teacher] 32 32 24.46 - 0.304 - - - -
5.1 Text-guided Image Generation
We also demonstrate effectiveness of DDIL with distribution matching techniques which adopt multi-step student like in DMD2. When we apply DMD2 to SSD1B, we can observe that DDIL improves FID from $31.77\to 27.72$ i.e., 12.7% improvement and clip score from $0.320\to 0.326$ and HPSv2 score from $0.302\to 0.304$.
Compared to 64 A100s with batchsize of 128 reported in DMD2, in our work we only use 7 A100 GPUs with batchsize of 7 and demonstrate comparable performance with DMD2 on SSD1B, even with worse critic especially when we adopt reflected diffusion to improve stability of training and then incorporating DDIL further improves performance. DDIL demonstrates strong performance with DMD2, LCM using significantly smaller batch sizes and fewer gradient steps demonstrating its generalization across methods and potential for wider appplication when distilling diffusion models.
Diversity vs. Quality trade-off: Diffusion distillation aims to compress a pre-trained diffusion model while maintaining performance. This compression is achieved by reducing compute i.e., iterations. However, this presents a fundamental challenge: achieving both high quality and diverse generations with significantly reduced model capacity. Within distillation our goal is to mimic trajectory or map of pretrained diffusion model in significantly fewer steps and hence this could result in discretization errors which compound across steps resulting in poor performance.
So depending on distillation method we choose fidelity or diversity, adversarial distillation methods [34, 35] exhibit a decrease in generation diversity compared to the baseline but demonstrate good quality. As discussed in previous section most common objectives in distribution matching adopt mode-seeking objectives [39, 44] where as methods like progressive distillation preserve map from prior to data-distribution and hence preserves diversity but looses on quality because of discretization errors, where as other methods can ignore high curvature or difficult regions of initial map there by choosing quality over diversity. There is need for future work to improve diversity, as both DMD2 and LCM have poor coverage, recall on SSD1B.
Table 3: Text guided image generation results on COCO 2014-30K validation set except for CLIP Score which is evaluated on COCO 2017-5K . These results are obtained by adopting latent consistency distillation retrained for SSD1B and incorporating DDIL within DMD2 setting. We integrate DDIL into the DMD2 framework by unrolling just the student model exactly like DMD2 but also unrolling teacher to corresponding noise level too to better capture underlying data distribution and align gradient fields of student model and teacher model.We use Guidance = 8 for teacher, LCM* checkpoint was not trained by us.
Model Steps FID@5K []delimited-[][$\downarrow$][ ↓ ] FID@30K []delimited-[][$\downarrow$][ ↓ ] CLIP []delimited-[][$\uparrow$][ ↑ ] HPSV2 []delimited-[][$\uparrow$][ ↑ ] Precision []delimited-[][$\uparrow$][ ↑ ] Recall []delimited-[][$\uparrow$][ ↑ ] Density []delimited-[][$\uparrow$][ ↑ ] Coverage []delimited-[][$\uparrow$][ ↑ ]
SSD1B (Teacher) 20 30.23 20.71 0.336 0.297 0.463 0.679 0.186 0.271
SSD1B-LCM* 4 35.23 28.46 0.311 0.282 0.336 0.586 0.117 0.171
SSD1B-DMD2 4 31.77 23.24 0.320 0.302 0.361 0.605 0.136 0.196
+DDIL 4 27.72 18.73 0.326 0.304 0.426 0.658 0.163 0.234
Improvement ($\mathbf{\Delta}$) 4 12.74% 19.41% 1.88% 0.66% 18.05% 8.76% 19.85% 19.38%
5.2 Covariate shift Analysis
We investigate covariate shift stemming from error accumulation in distilled diffusion models. Our analysis utilizes a 32-step Classifier-Free Guidance (CFG) teacher and a 4-step student trained via progressive distillation (PD) and evaluate on the MS-COCO 2017 (5K) validation set. We conduct mixed-rollout evaluations, leveraging a stateless DDIM solver to stochastically select between the teacher and student at each step (t-250), with a prior probability $p_{T}$ of choosing the teacher.
Given that we are working with PD checkpoint, if there is no discrepancy between teacher (black) trajectory and student’s trajectory (red+blue) in Fig. 3 we should observe same performance irrespective of relative prior on teacher, i.e., even when we switch to teacher on student’s trajectory (green transitions) in Fig. 3 performance should not improve. In Tab. 9 we control choosing teacher at different intermediate time steps and in absence of covariate shift increasing teacher selection rate (green transitions) should not vary performance but we can see that as we increase prior on teacher we see consistent improvement in CLIP score and FID validating hypothesis on presence of covariate shift.
Table 4: Covariate shift Analysis: Evaluation with different teacher selection rates.
$p_{T}$ FID [$\downarrow$] CLIP [$\uparrow$]
0.8 23.14 0.319
0.6 22.33 0.317
0.4 21.95 0.313
0.2 21.92 0.307
0.0 22.42 0.302
6 Conclusion
This work introduces DDIL, a novel framework for distilling diffusion models that addresses the challenge of covariate shift while preserving the marginal data distribution. Integrating DDIL with established distillation techniques like LCM, DMD2 and PD consistently yields quantitative and qualitative improvement without loss of diversity demonstrating generalization across various methods. This is particularly noteworthy given the already high performance of distilled models, which often rival their teacher models with limited headroom for improvement.
Furthermore, we also show that integrating DDIL within the DMD2 framework enhances training stability, achieves better diversity than the SSD1B teacher model, even with a significantly smaller batch size (7 compared to 128 in the original DMD2 training setup). This highlights computational efficiency, diversity preserving and practical usefulness of DDIL across distillation methods.
Appendix A Additional Training Details
In case of progressive distillation (PD), we use DPM++ multi-step [21] solver (2-step) for all evaluations except co-variate shift analysis, where a state-less solver like DDIM [45] will enable switching between two different models and corresponding reverse processes. We train our model on an internal text-to-image dataset and only perform distillation without updating the teacher model. In case of consistency distillation with LCM, we adopt the default LCMScheduler. In the case of DMD2, we follow the same settings as the original repository except that in our work, we focus on SSD1B [6] instead of SDXL to reduce computational requirements for training. By default, we adopt reflected diffusion for all three diffusion distillation techniques we considered in this work.
Distribution Matching Distillation(DMD2):
In case of DMD2, to integrate DDIL in addition to unrolling student model we also unroll pre-trained or teacher model. This makes student model to be well aligned with pre-trained diffusion model and also better capture underlying statistics. In addition this obtains better feedback within score distillation framework as we not only obtain teacher’s feedback w.r.t student’s generation at terminal states but w.r.t each of student’s denoising step when input to student model is initialized by unrolling pre-trained teacher model until corresponding input timestep.
When integrating DDIL within DMD2, we do not consider forward diffusion and instead unroll teacher model and assume pre-trained diffusion model is a proxy for marginal data distribution. In principle we can consider real data to better capture data distribution i.e., tighter ELBO but we leave it for future work. To integrate DDIL within DMD2, we found sampling priors of $\beta_{teach\_bckwrd}=0.6,\beta_{student\_bckwrd}=0.4$ and as we don’t consider real data to obtain score distillation feedback $\beta_{frwd}=0.0$
Latency Consistency Models(LCM):
In case of LCM [22], we explored a setting where we add backward trajectories while distillation in addition to forward diffusion, i.e., we found significant boost even without unrolling teacher. This would be mean more sampling at training compared to interactive denoising considered in case of Progressive Distillation, where we collect an independent dataset exactly following DAgger [30] protocol. We can consider either protocols to incorporate DDIL with distillation methods. In case of LCM 50% of time we train following standard protocol and remaining 50% of time, we unroll student model with LCMScheduler to a random step within total number of inference steps $\in{2,3,4,5,6}$ to obtain feedback to correct covariate shift within consistency distillation formulation. We adopt sampling priors of $\beta_{frwd}=0.5,\beta_{teach\_bckwrd}=0.0,\beta_{student\_bckwrd}=0.5$ in case of LCM.
Progressive Distillation:
For $16\to 8$, we follow a similar protocol as $32\to 16$, but we further split each stage of training into two parts. First, we do distillation for 6K steps to obtain a checkpoint and resume with warmup and $051e-$ 05 learning rate for another 4K steps of training. For $8\to 4$ and $4\to 2$ distillation, we first distill model for 8K steps followed by another 6K steps.We split single stage of training into two parts to exactly match training protocol of Step Distillation and DDIL. See supplemental for additional details. We adopt timesteps for discretization from the default config of the DPM++ solver i.e., for 4-step models our timesteps are $\{999,749,500,250\}$.
We follow similar protocol as progressive distillation for PD+DDIL but as we now distill using on text-to-image data using forward process and interactively mixed unrolled trajectories, we have this additional sampling prior hyper parameters $\beta_{frwd},\beta_{teach\_bckwrd},\beta_{student\_bckwrd}$.
Sampling Priors for PD: For $16\to 8$ distillation $75\%$ we obtain $z_{t}$ using forward diffusion i.e, $\beta_{frwd}=0.75$ making this part of training exactly equivalent to step Distillation. For remaining $25\%$ we choose data from mixed interaction unrolled trajectories where for first 6K steps, student is only selected $15\%$ of time for $z_{t}\to z_{t-2k}$ transition for trajectory collection, where as for next 4K iterations we choose student $80\%$ of time and train $50\%$ of time on collected trajectories.
Effectively for first 6K steps of training we have prior $x\sim p_{data}(x)$ as $\beta_{frwd}=0.75$, $\beta_{teach\_bckwrd}=0.25*0.85=0.2125$ and $\beta_{student\_bckwrd}=0.25*0.15=0.0375$. And for next 4K iterations, $\beta_{frwd}=0.5$ and $\beta_{teach\_bckwrd}=0.5*0.2=0.10$ and $\beta_{student\_bckwrd}=0.5*0.8=0.40$.
And follow same sampling prior protocol for two parts of training for each stage of progressive distillation with same hyper-parameters as step distillation(PD). Overall we need ${}40K$ updates to obtain a 4-step checkpoint or ${}55K$ updates to obtain 2-step checkpoint for PD or PD+DDIL, as our sampling is parallelized across GPUs we observe $<5\%$ overhead for DDIL over step distillation, it takes 2 days on single node of 8 A100 GPUs to perform distillation.
Appendix B Additional Evaluation Results
@@
SnapFusion* [15] 8 16 24.20 0.300 -
Step Distillation* [25] 8 8 26.90 0.300 -
Step Distillation* [25] 4 4 26.40 0.300 -
UFOGen* [41] 1 1 22.5 0.311 -
ImagineFlash* $\ddagger$ [12] 2 2 34.7 0.301 -
@@
+ Reflected 4 4 24.25 (-0.6%) 0.306 0.59
+ DDIL 4 4 23.44(-3.9%) 0.308 0.59
+ Reflected + DDIL 4 4 22.86 (-6.3%) 0.309 0.59